find the logarithm of the following
(i) 23.82
(ii) 0.006735
find the anti-logarithm of the following
(i) 0.8401
(ii) 2.8370

Answer by lwsshak3(11628)

find the logarithm of the following
(i) 23.82
(ii) 0.006735
find the anti-logarithm of the following
(i) 0.8401
(ii) 2.8370
**
Using a calculator, assuming base 10.
(i) log 23.82=1.3769..
(ii)log 0.006735 =-2.1717..
find the anti-logarithm of the following
(i)anti log 0.8401=10^(.8401)=6.9199..
(ii)antil og 2.8370=10^(2.8370)=687.0684..